OpenStudy (loser66):

Let m be an integer, \(m\geq 2\), and let \(Z_m \)be the set of all positive integers less than m, \(Z_m = {0,1,.....,m-1}\) If \(\alpha ~~and~~\beta\) are in \(Z_m\), let \(\alpha +\beta\)be the least positive remainder obtained by dividing the sum of \(\alpha ~~and~~\beta\) by m Similarly, let \(\alpha\beta\) be the least positive remainder obtained by dividing the product of \(\alpha ~~and~~\beta\) by m. Prove: a) \(Z_m\) is a field if and only if m is a prime b) What is -1 in \(Z_5\) ? c)What is 1/3 in \(Z_7\) ? Please, help

OpenStudy (anonymous):

i'm not your huckleberry on this stuff. try sattelite

OpenStudy (anonymous):

b) it's 4. \[4 \times 1 \equiv 4\equiv -1 \mod 5\]\[4 \times 2 \equiv 3\equiv -2 \mod 5\]\[4 \times 3 \equiv 2\equiv -3 \mod 5\]\[4 \times 4 \equiv 1\equiv -4 \mod 5\]

OpenStudy (loser66):

I got this , friend. In general, it is a multiple of 4, right?

OpenStudy (cail):

i dont know this, sorrry. Thanks for helpinggg

OpenStudy (anonymous):

c. 3x+7y=1 => 3(5)+7(-2)=1 => \(3\cdot5\equiv 1 \mod 7\), thus multiplying by 5 is equivalent to dividing by 3, mod 7. Ex. \( 3\cdot 5\equiv15\equiv1\mod 7 \) and \(3\div 3=1\) \( 6\cdot 5\equiv30\equiv2\mod 7 \) and \(6\div 3=2\) \( 9\cdot 5\equiv45\equiv3\mod 7 \) and \(9\div 3=3\) \( 12\cdot 5\equiv60\equiv 4\mod 7 \) and \(12\div 3=4\) and so on...

OpenStudy (loser66):

I am reading, not digest yet, hihihi... spending time to help other, that's why. I am sorry for that

OpenStudy (loser66):

I have a question there: 3*5 =1 (mod7) but 7*(anything) = 0 (mod 7) why do you choose -2?

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

Beats me.

OpenStudy (loser66):

a)

OpenStudy (loser66):

for c) I was taught the remainder in Z_7 is an integer. I am not sure how to get 1/3

OpenStudy (anonymous):

Let \(n=1/3\) that means \(3n=1\) and we know the definition for multiplication.

OpenStudy (loser66):

got you,

OpenStudy (anonymous):

For a, you probably need to use the division algorithm \[ n = dq+r \]

OpenStudy (anonymous):

what can you use for your proof?

OpenStudy (loser66):

I don't know, it's my lecture for next semester, my prof asks us preview before the course. I read the book and try some exercises

OpenStudy (anonymous):

oh sorry i came late you are figuring out what \(\frac{1}{3}\) is in \(\mathbb{Z}_7\)

OpenStudy (anonymous):

i have never seen it written like this before, usually it is written \(3^{-1}\) i.e. the multiplicative inverse of 3 there are no fractions in \(\mathbb{Z}_7\)

OpenStudy (anonymous):

since you only have 7 elements, the best method is to grind it till you find it

OpenStudy (anonymous):

\(1\times 3=3\) \[2\times 3=6\]\[3\times 3=2\]\[4\times 3=5\]\[5\times 3=1\] got it on the fifth try

OpenStudy (loser66):

you said that there is no fraction in Z_7??

OpenStudy (anonymous):

yes, that is what i said, and i think i will stick by it

OpenStudy (anonymous):

but i guess it is just a matter of notation is all since \(\mathbb{Z}_7\) is a field every non zero element has a multiplicative inverse but it is usually denoted as \(a^{-1}\) not \(\frac{1}{a}\)

OpenStudy (loser66):

so, it can be the answer, right? we can have "NO SOLUTION" as a solution, right?

OpenStudy (anonymous):

no no, the answer is 5

OpenStudy (anonymous):

\(5\times 3=1\) so \(5=3^{-1}\)

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio Let \(n=1/3\) that means \(3n=1\) and we know the definition for multiplication. \(\color{blue}{\text{End of Quote}}\) Definition of multiplication is used to evaluate division.

OpenStudy (anonymous):

what @wio said

OpenStudy (anonymous):

just never seen it written as \(\frac{1}{3}\) before is all

OpenStudy (anonymous):

\(1/3\) is not a number, but an algebraic expression to be parsed.

OpenStudy (anonymous):

kk got it, just never saw it

OpenStudy (anonymous):

For the proof, you probably want to factor \(m\) into its prime decomposition and consider the consequences.

OpenStudy (anonymous):

Proving that something is a field can be pretty tedious, though.

OpenStudy (loser66):

oh, I used to learn this concept from my discrete math 2 years ago but just a little bit.

OpenStudy (anonymous):

snap proof, in \(\mathbb{Z}\) the ideal \(p\mathbb{Z}\) is maximal, making \(\mathbb{Z}_p\cong \mathbb{Z}/p\mathbb{Z}\) a field but i have a feeling that is not what you are being asked to do what you have to show is that any \(n\in\mathbb{Z}_p\) has a multiplicative inverse

OpenStudy (cail):

can you please help me @satellite73

OpenStudy (anonymous):

all the other properties of being a field are routine, this is the only one that needs work, and i would suggest "bezout" or whatever it is called

OpenStudy (loser66):

@satellite73 I know Bezout theorem

OpenStudy (anonymous):

proof is something like if \(m\in \mathbb{Z}_p\) then since \(p\) is prime, \(m\) and \(p\) are relatively prime therefore there are \(x,y\) with \(1=xm+yp\) this gives \[xm=1-yp\] etc

OpenStudy (anonymous):

there is more to it, but that is the basic idea all other field properties are inherited from the regular properties of integers

OpenStudy (anonymous):

yeah bezout. same thing as finding a unit fraction mod n. ax+by=1 => ax = 1 mod y => a = 1/x mod y that's all i did. you were specifically looking for the inverse of 3 so x = 3.

OpenStudy (loser66):

Thanks all, now I know where I can start learning.